Altamonte Telecommunications has a target capital structure that consists of 40% debt and 60% equity. The company anticipates that its capital budget for the upcoming year will be $1,000,000. If Altamonte reports net income of $2,500,000 and it follows a residual dividend payout policy, what will be its dividend payout ratio? Round your answer to two decimal places.

  %
Answer #1

Dividend payout=Net income-(Capital budget*Weight of equity)

=2,500,000-(0.6*1,000,000)=1,900,000

Dividend payout ratio=Dividend payout/Net income

=1,900,000/2,500,000

=76%